How many molecules and atoms of sulphur are present in 0.1 mole of S8 molecules?

Grade:10

1 Answers

AVNISH SINGH
6 years ago
1 mole = 6.022*1023 molecules
0.1 moles =0.1*6.022*1023
                   = 6.022*1022 molecules
1 molecule of sulphur contain 8 atoms 
no. of atoms = 8* no. of molecules
                    =4.8176*1023 atoms